OpenStudy (anonymous):

Using the distributive property, which number completes the statement correctly? –4 • (–6 + 5.7) = –4 • (–6) +_____?• 5.7

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

@dan815 @Whitemonsterbunny17

OpenStudy (anonymous):

\[a \left( b+c \right)=ab+ac\]

OpenStudy (anonymous):

is it -4?

jimthompson5910 (jim_thompson5910):

Yes it is because you multiply the outer term (-4) by each of the inner terms (-6 and 5.7). Then you add up those products.

OpenStudy (calculusxy):

-4

OpenStudy (anonymous):

Thank you @jim_thompson5910